RELATIVISTIC JAHN-TELLER EFFECTS IN THE QUARTET STATES OF K_{3} AND RB_{3}: A VIBRA-TIONAL ANALYSIS OF THE 2^{4}E′ ← 1^{4}A′_{2} ELECTRONIC TRANSITIONS BASED ON AB INITIO CAL-CULATIONS

\underline{A. W. HAUSER}, G. AUBÖCK, C. CALLEGARI and W. E. ERNST, Institute of Experimental Physics, Graz University of Technology, Petersgasse 16, A-8010 Graz, Austria.

We apply the Multireference Rayleigh Schrödinger Perturbation Theory of second order to obtain the adiabatic potential energy surface of the 1^{4}A′_{2} electronic groundstate and the 2^{4}E′ excited state of K_{3} and Rb_{3}. Both trimers show a typical E×e Jahn-Teller distortion in their 2^{4}E′ state, which is analyzed in terms of the relativistic Jahn-Teller effect theory. Linear, quadratic as well as spin-orbit coupling terms are extracted from the ab initio results and used to obtain theoretical spectra for a direct comparison to laser-induced fluorescence and magnetic circular dichroism spectra of alkali-doped helium nanodroplets [Auböck et al. J. Chem Phys. 129 114501 (2008)].